# Large mechanical squeezing through synthetic magnetism in optomechanics

### v1

@@ IV Conclusion
We have investigated on a synthetic magnetism effect inducing mechanical squeezing enhancement in an optomechanical system, which hosts a backward stimulated Brillouin scattering process. Our benckmark system consists of an acoustic mode that couples to two optical modes through the BSBS process, and a nonlinear mechanical oscillator that couples to the two optical modes through the standard optomechanical radiation pressure. A mechanical coupling, with a strengh $J_{m}$ that is modulated through a phase $\theta$, is connecting the acoustic and mechanical modes, which induces a synthetic magnetism in our proposal. Without this synthetic magnetism, there is a given amount of squeezing that is generated in the system. This squeezing is mainly induced by the BSBS process, and is fragile against thermal noise. When the synthetic magnetism is accounted, the degree of the generated squeezing is greatly enhanced and goes far beyond the 3dB. Moreover, this large induced squeezing persists even when the system is free from the BSBS process. Furthermore, this generated squeezing is robust enough against thermal noise compared to the case without synthetic magnetism. Another merit of the synthetic magnetism in our proposal is revealed through the peaks and dips of both squeezing magnitude and mechanical effective phonon number depending on the modulation phase of the mechanical coupling. This oscillatory feature is reminscent of a sudden death and revival of squeezing phenomenon, which can be used to maintain a desired magnitude of squeezing by tuning the phase $\theta$. Our work sheds light on a flexible scheme that can be used to generate a large amount of mechanical squeezing, far beyond the $\rm{3dB}$ limit. Our scheme can be implemented in optical and microwaves cavities, as well as in hybrid optomechanical systems. Such a generated squeezed states can be useful for range of quantum applications including quantum information processing, quantum sensing/metrology, and the recent development in quantum computing.
@@ Acknowledgments
Funding: This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa. P. Djorwe acknowledges the receipt of a grant from the APS-EPS-FECS-ICTP Travel Award Fellowship Programme (ATAP), Trieste, Italy. S. Agasti wishes to acknowledge the European Union; Project number: 101065991 (acronym: SingletSQL) for supporting the work. A.K. Sarma acknowledges the STARS scheme, MoE, government of India (Proposal ID 2023-0161).
Appendix A Linearization
Appendix B Bogoliubov transformation

### v2

@@ IV Conclusion
We have investigated the synthetic magnetism effect inducing mechanical squeezing enhancement in an optomechanical system, which hosts a backward stimulated Brillouin scattering process. Our model system consists of an acoustic mode that couples to two optical modes through the BSBS process, and a nonlinear mechanical oscillator that couples to the two optical modes through the standard optomechanical radiation pressure. A mechanical coupling, with a strength $J_{m}$ that is modulated through a phase $\theta$, is connecting the acoustic and mechanical modes, which induces a synthetic magnetism in our proposal. Without this synthetic magnetism, there is a given amount of squeezing that is generated in the system. This squeezing is mainly induced by the BSBS process, and is fragile against thermal noise. When the synthetic magnetism is accounted, the degree of the generated squeezing is greatly enhanced and goes far beyond the 3dB. Moreover, this large induced squeezing persists even when the system is free from the BSBS process. Furthermore, this generated squeezing is robust enough against thermal noise compared to the case without synthetic magnetism. Another merit of the synthetic magnetism in our proposal is revealed through the peaks and dips of both squeezing magnitude and mechanical effective phonon number depending on the modulation phase of the mechanical coupling. This oscillatory feature is reminiscent of a sudden death and revival of squeezing phenomenon, which can be used to maintain a desired magnitude of squeezing by tuning the phase $\theta$. Our work sheds light on a flexible scheme that can be used to generate a large amount of mechanical squeezing, far beyond the $\rm{3dB}$ limit. Our scheme can be implemented in optical and microwaves cavities, as well as in hybrid optomechanical systems. Such a generated squeezed states can be useful for a range of quantum applications including quantum information processing, quantum sensing and metrology, and the recent development in quantum computing.
@@ Acknowledgments
This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa. S. Agasti wishes to acknowledge the European Union H2020 MSCA; Project number: 101065991 (acronym: SingletSQL) for supporting the work. A.K. Sarma acknowledges the STARS scheme, MoE, government of India (Proposal ID 2023-0161). K.S. Nisar is grateful to the funding from Prince Sattam bin Abdulaziz University, Saudi Arabia project number (PSAU/2024/R/1445). The authors are thankful to the Deanship of Graduate Studies and Scientific Research at University of Bisha for supporting this work through the Fast-Track Research Support Program.
Appendix A Effective Hamiltonian
The Hamiltonian that describes the system considered by us is given by,
$H$ $=\sum_{j=1,2}\omega_{c_{j}}a_{j}^{\dagger}a_{j}+\omega_{a}b_{a}^{\dagger}b_{a}+\omega_{m}b_{m}^{\dagger}b_{m}-\sum_{j=1,2}g_{m,a}a_{j}^{\dagger}a_{j}(b_{m}+b_{m}^{\dagger})$ (20)
$+\frac{\eta}{2}(b_{m}+b_{m}^{\dagger})^{4}+J_{m}({e^{i\theta}}{b_{a}^{\dagger}}{b_{m}}+{e^{-i\theta}}{b_{a}}{b_{m}^{\dagger}})$
$+\sum_{j=1,2}iE_{j}(a_{j}^{\dagger}e^{-i\omega_{p_{j}}t}-a_{j}e^{i\omega_{p_{j}}t})-g_{a}(a_{1}^{\dagger}a_{2}b_{a}+a_{1}a_{2}^{\dagger}b_{a}^{\dagger}).$
The total Hamiltonian of the hybrid system in a frame rotating with the laser frequency $\omega_{p_{j}}$ is given by:
$H^{\prime}$ $=-\sum_{j=1,2}\Delta_{j}a_{j}^{\dagger}a_{j}+\Delta_{a}b_{a}^{\dagger}b_{a}+\omega_{m}b_{m}^{\dagger}b_{m}-\sum_{j=1,2}g_{m,a}a_{j}^{\dagger}a_{j}(b_{m}+b_{m}^{\dagger})$ (21)
$+\frac{\eta}{2}(b_{m}+b_{m}^{\dagger})^{4}+J_{m}({e^{i\theta}}{b_{a}^{\dagger}}{b_{m}}+{e^{-i\theta}}{b_{a}}{b_{m}^{\dagger}})$
$+iE_{1}(a_{1}^{\dagger}-a_{1})+iE_{2}(a_{2}^{\dagger}-a_{2})-g_{a}(a_{1}^{\dagger}a_{2}b_{a}+a_{1}a_{2}^{\dagger}b_{a}^{\dagger}),$
with $\Delta_{j}=\omega_{p_{j}}-\omega_{c_{j}}$, and $\Delta_{a}=\omega_{a}+\omega_{p_{2}}-\omega_{p_{1}}$. By considering that the control field $a_{2}$ is strong enough compared to $a_{1}$, it can be treated classically by deriving its steady-state as,
$\dot{a}_{2}$ $=i[H^{\prime},a_{2}],$ (22)
$=(i\Delta_{2}-\frac{\kappa_{2}}{2})a_{2}+ig_{m}a_{2}(b_{m}+b_{m}^{\dagger})+E_{2}+ig_{a}a_{1}b_{a}^{\dagger}$
$=(i\Delta^{\prime}_{2}-\frac{\kappa_{2}}{2})a_{2}+E_{2}+ig_{a}a_{1}b_{a}^{\dagger},$ (23)
with $\Delta^{\prime}_{2}=\Delta_{2}+g_{m}a_{2}(b_{m}+b_{m}^{\dagger})$. The steady-state solution ($\dot{a}_{2}=0$) yields,
$\alpha_{2}$ $\sim\frac{-E_{2}}{i\Delta^{\prime}_{2}-\frac{\kappa_{2}}{2}}\hskip 10.00002pt\text{or}\hskip 10.00002pt|\alpha_{2}|\sim\frac{E_{2}}{\sqrt{\Delta_{2}^{{}^{\prime 2}}+\frac{\kappa_{2}^{2}}{4}}}.$ (24)
By substituting this expression in the rest of the Hamiltonian in Equation 21, we get the following reduced Hamiltonian,
$H^{\prime}$ $=-\Delta_{1}a_{1}^{\dagger}a_{1}+\Delta_{a}b_{a}^{\dagger}b_{a}+\omega_{m}b_{m}^{\dagger}b_{m}-g_{m}a_{1}^{\dagger}a_{1}(b_{m}+b_{m}^{\dagger})$
$+\frac{\eta}{2}(b_{m}+b_{m}^{\dagger})^{4}+J_{m}({e^{i\theta}}{b_{a}^{\dagger}}{b_{m}}+{e^{-i\theta}}{b_{a}}{b_{m}^{\dagger}})$ (25)
$+iE_{1}(a_{1}^{\dagger}-a_{1})-{G_{a}}(a_{1}^{\dagger}b_{a}+a_{1}b_{a}^{\dagger}),$
where the acoustic effective coupling is $G_{a}=g_{a}\alpha_{2}$ as mentionned in the main text.
Appendix B Oscillatory behavior and effect of the effective couplings
This section highlights some phenomena aforementioned in the main text related to Figure 3 and Figure 4. Indeed, it was mentioned that the position variance and the phonon number displayed in Figure 3 exhibit sort of oscillatory behavior depending on the synthetic phase $\theta$ around the phonon hopping rate $J_{m}=0.1\omega_{m}$. Moreover, it was pointed out that these two quantities are out of phase with each other over $\theta$. Figure 6a depicts the position variance while Figure 6b shows variation of effective phonon numbers against phase $\theta$. These depictions are extracted from Figure 3a and Figure 3b at $J_{m}=0.1\omega_{m}$, respectively. As expected, it can be clearly seen how these two quantities exhibit oscillatory features and are out of phase.
Figure 6: Mechanical position variance (a) and effective phonon number (b) (extracted from Figure 3 at $J_{m}=0.1\omega_{m}$) against variation of phase $\theta$. The other parameters are the same as in Figure 2.
In Figure 4, it has been pointed out that as the effective coupling $G_{a}$ and $G_{m}$ are enhanced, the strength of squeezing gets enhanced as well. To get a clearer view of that feature, Figure 7 displays the mechanical position variance in $G_{a}$ and $G_{m}$ space for $J_{m}=0$ (a) and $J_{m}=0.1\omega_{m}$ (b). As it can be seen in both figures, the squeezing get stronger and stronger as the strength of the two effective couplings increases. This justifies our statement aforementioned in the main text. Moreover, by paying attention to the colorbars in Figure 7, it appears that the squeezing is enhanced as the synthetic magnetism is accounted in the system, i.e., $J_{m}\neq 0$ and $\theta=\pi/2$. This can be seen by comparing the couple of points highlighted in Figure 7 for instance.
Figure 7: Squeezing degree of the mechanical position variance in $G_{a}$ and $G_{m}$ space for $J_{m}=0$ (a) and $J_{m}=0.1\omega_{m}$ (b). The rest of the parameters are the same used in Figure 2.